4.     A coyote jumps over the edge of a cliff with a speed of 20 m/s. It hits the surface of the canyon 15 s later.


(a) Calculate the vertical speed of the coyote when it reaches the canyon surface.

(b) How high is the cliff?

(c) How far from the foot of the cliff will the coyote land?


5.     An aircraft flying horizontally at 150 m/s, drops a bomb which hits the target after 8 s.

Find:

a)   the distance travelled horizontally by the bomb

b)   the vertical speed of the bomb at impact

c)   the distance travelled horizontally by the aircraft as the bomb fell


1
Expert's answer
2021-04-20T16:49:35-0400

"\text {body formulas thrown horizontally:}"

"V_x = const ; S_x=V_x*t"

"V_y =g*t;S_y=\frac{gt^2}{2}"


"4. V_x= 20m\/s; t =15 s"

"a)V_y =g*t=9.8*15=147m\/s"

"b)S_y=\frac{gt^2}{2}=\frac{9.8*15^2}{2}=1102.5m"

"c) S_x=V_x*t = 20*15 = 300m"


"5.V_x = 150m\/s ;t = 8s"

"a) S_x=V_x*t = 150*8 = 1200m"

"b)V_y =g*t=9.8*8=78.4m\/s"

"c) \text{the speed of the plane and the horizontal speed }\newline
\text{of the bomb are the same hence }S=S_x=1200m"


Answer:

4.a)147 m/s b)1102.5 m c)300 m

5.a)1200 m b) 78.4 m/s c) 1200 m